The owner of Boston sports bar believes that, on average, her restaurant is busier on days when the Red Sox play an away game (a game played at another team’s stadium), but she wants to be sure before adding more staff. To test whether this is true, she takes a random sample of 50 days over the course of the baseball season and records the total daily revenue, along with whether the Red Sox were playing away that day (1 if yes, 0 if no). Using the data provided, perform a regression analysis to determine the effect of Red Sox away games on revenue. Be sure to include the residuals and residual plot in your analysis.

USE EXCEL PLEASE5 Kezdőlap chegg 1 - Excel (A termékaktiválás nem sikerült) Nézet Mondja el, mit szeretne tenni..., Fail Beszúrás Lapelrendez5 Kezdőlap chegg 1 - Excel (A termékaktiválás nem sikerült) Nézet Mondja el, mit szeretne tenni..., Fail Beszúrás Lapelrendez
Answer #1

Answer: ---- The regression equation is defined as, Date: ---28/4/2020 Y = Bo + B1X Where, dependent variable Y is the DailyOME INSERT PAGE LAYOUT FORMULAS A ZA EH ! Data Analysis DATA REVIE Data Analysis Analysis Tools OK Cancel Help Histogram MoviDaily Red Sox away 1 Revenue ($) games 1560 2918 Regression Input Input Y Range: ок SA$1:$A$51 Cancel Input X Range: $B$1:$B$

B C D E F SUMMARY OUTPUT Regression Statistics 4 Multiple R 0.474566 5 R Square 0.225213 6 Adjusted R Square 0.209072 7 Stand

RESIDUAL OUTPUT Observation | Predicted Daily Revenue (5) Residuals 2264.5714 -704.5714 2264.5714 653.4286 2264.5714 276.4286

1768.3182 429.6818 2264.5714 -294.5714 1768.3182 602.6818 2264.5714 -589.5714 1768.3182 396.6818 1768.3182 -755.3182 2264.571

1768.3182 29.6818 2264.5714 -90.5714 2264.5714 478.4286 1768.3182 410.6818 2264.5714 -118.5714 2264.5714 -312.5714 1768.3182

2264.5714 690.4286 2264.5714 100.4286 2264.5714 -584.5714 1768.3182 669.6818 46 47 1768.3182 -400.3182 2264.5714 -352.5714 17

Red Sox away games Residual Plot 1000 Residuals 0 0.2 0.4 0.6 0.8 -5003 -1000 Red Sox away games from the regression output s

When Red Sox were not playing away (X = 0) Y = 1768.3182 + 496.2532 x 0 = 1768.3182 Daily revenue is $ 1768.3182 There is dif